6) 1413 + 92n2 – 421 8) 962 – 33b + 30 MacBGAS

Factor the equation
Answer #1

лч ть+кин) и 2 (17 - 91 +, + 2 чі - ( – at - h FчE) из - в (+/n+ 1) = (n + 1) = | 24 (n+1) -32 4. чь — 33 + 30 — 21з°— ть+ (